OpenStudy (anonymous):

lim-> 0(pi/2) 90sin0

geerky42 (geerky42):

\[\large \lim_{\theta\to\frac\pi2} ~90\sin\theta\]?

geerky42 (geerky42):

Just do direct substitution.

OpenStudy (anonymous):

Not sure how to do that. Thank you for responding.